Multiply and simplify. Assume variables represent nonzero real numbers.
c^15*c^0
: Multiply and simplify. Assume variables represent nonzero real numbers.
c^15*c^0

Answer by checkley77(3848)

c^15c^0=c^15*1=c^15 answer.